OpenStudy (anonymous):

Solve 3x − x + 2 = 12. −5 4 7 5

OpenStudy (texaschic101):

3x - x + 2 = 12 --- subtract 12 from both sides 3x - x = 12 - 2 -- simplify 2x = 10 -- divide both sides by 2 x = 5

OpenStudy (texaschic101):

oops...I meant subtract 2 from each side

OpenStudy (anonymous):

Thankyou so much !!!!

OpenStudy (texaschic101):

anytime :)